### v1

# Dimension of Inhomogeneous Sub-Self-Similar Sets

Abstract.
In this paper, we introduce the concept of Inhomogeneous sub-self-similar (ISSS) sets, building upon the foundations laid by Falconer (Trans. Amer. Math. Soc. 347 (1995) 3121–3129) in the study of sub-self-similar sets and drawing inspiration from Barnsley’s work on inhomogeneous self-similar sets (Proc. Roy. Soc. London Ser. A 399 (1985), no. 1817, 24). We explore a range of examples of ISSS sets and elucidate a method to construct ISSS sets. We also investigate the upper and lower box dimensions of ISSS sets and discuss the continuity of the Hausdorff dimension.
Key words and phrases:
Self-similar sets, Inhomogeneous IFS, Hausdorff dimension, Box dimension, Open-set condition
Key words and phrases:
Self-similar sets, Inhomogeneous IFS, Hausdorff dimension, Box dimension, Open-set condition
2010 Mathematics Subject Classification:
Primary 28A80; Secondary 28A78, 26A18
1. Introduction
A wide variety of fractals exhibit self-similarity, i.e., they consist of infinitely smaller replicas that resemble the whole. Hutchinson [20] provided a comprehensive framework for self-similar sets and constructed them mathematically using Iterated Function Systems (IFS). The construction of self-similar sets through IFS has been extensively studied and popularized by Barnsley and his collaborators [3, 5, 4]. Barnsley [3] introduced the concept of fractal interpolation functions (FIF), and many researchers have since worked on FIFs and their applications. For more details, interested readers may see [8, 7, 23, 26, 32, 35, 37, 36, 21]. Barnsley and Demko [4] extended the IFS framework by introducing inhomogeneous IFS, which includes traditional IFS along with a compact set called a condensation set. Inhomogeneous IFS generalizes the concept of IFS in such a way that, by considering the condensation set as the empty set, it reduces to the classical self-similar case. Furthermore, Mauldin and Williams [30] introduced the graph-directed IFS (GD-IFS), which generalizes traditional IFS. In this regard, Boore and Falconer [6] provide an example of a GD fractal that cannot be generated by a traditional IFS. Recently, Dubey and Verma further generalized GD-IFS by introducing the concept of inhomogeneous GD-IFS (see [10, 11, 33] for more details). In [14], Falconer introduced the concept of sub-self-similar (SSS) sets by replacing equality in the self-similar equation with set inclusion. Computing the exact dimension of a fractal set remains a challenging task in fractal geometry. Various methods exist for estimating the dimensions of fractals. Chandra and Abbas [9] computed the fractal dimension of mixed Riemann–Liouville integrals using the $\delta$-covering method. Falconer [12] discussed fractal dimensions using the Fourier transform method. Priyadarshi [34] estimated the lower bound on the Hausdorff dimension of sets of complex continued fractions using operator-theoretic techniques. Chandra and Abbas [8] studied the fractal dimension of fractal functions using function space techniques. Falconer [12] also explored fractal dimensions using potential-theoretic techniques and the mass distribution principle. Hochman [19] investigated the Hausdorff dimension of self-similar sets and measures on the real line using entropy techniques and provided an inverse theorem for the entropy of convolution. Shmerkin [40] extended this work by studying the $L^{q}$-dimension of dynamically driven self-similar measures. Following Falconer’s work [14] and using the concept of inhomogeneous self-similar sets, we introduce and construct inhomogeneous sub-self-similar sets (ISSS). Motivated by Fraser’s work [15], we estimate the fractal dimension of these sets using the $\delta$-covering method.
The outline of the paper is as follows: Section 2 provides the preliminaries. In Section 3, we present examples of ISSS sets. Section 4 discusses one of the methods for constructing ISSS sets, while Sections 5 and 6 are devoted to estimating the fractal dimension of these sets and discussing the continuity of the Hausdorff dimension for ISSS sets. In Section 7, we study the self-similar structure on the product of inhomogeneous IFS. Finally, Section 8 concludes the paper with remarks and outlines potential future directions for research.
2. Preliminaries
In this section, we recall the notion of sub-self-similar sets and introduce the concept of inhomogeneous sub-self-similar sets.
Let $(X,d)$ be a complete metric space, and let $f_{1},f_{2},\dots,f_{N}$ be similarity mappings on $X$. A unique non-empty compact set $A$ is called the self-similar set associated with the iterated function system (IFS) $\mathcal{I}:=\{X;f_{1},f_{2},\dots,f_{N}\}$ if $A$ satisfies the following equation, known as the self-similar equation:
$A=\bigcup_{i=1}^{N}f_{i}(A).$ (1)
Falconer [14] generalized the concept of self-similar sets by replacing equality in the self-similar equation with set inclusion. A non-empty compact set $F\subseteq X$ is called a sub-self-similar set (SSS set) if it satisfies:
$F\subseteq\bigcup_{i=1}^{N}f_{i}(F).$ (2)
Given a compact set $C\subseteq X$, referred to as the condensation set, a unique non-empty compact set $A_{C}$ is called the inhomogeneous self-similar set associated with the IFS $\mathcal{I}:=\{X;f_{1},f_{2},\dots,f_{N},C\}$ if it satisfies the following equation, known as the inhomogeneous self-similar equation:
$A_{C}=\bigcup_{i=1}^{N}f_{i}(A_{C})\cup C.$ (3)
Analogously, we define inhomogeneous sub-self-similar sets (ISSS sets), which extend the concept of inhomogeneous self-similar sets by replacing equality in (3) with set inclusion. A non-empty compact set $F_{C}\subseteq X$ is called an ISSS set if it satisfies:
$F_{C}\subseteq\bigcup_{i=1}^{N}f_{i}(F_{C})\cup C.$ (4)
Recall that if $N_{\delta}(F)$ represents the minimum number of sets of diameter $\delta$ covering $F$, then the upper and lower box-dimension of a non-empty compact set $F\subset\mathbb{R}^{n}$ is given by
$\overline{\dim}_{B}(F)=\limsup_{\delta\to 0^{+}}\frac{\log N_{\delta}(F)}{-\log\delta},~\text{and}~\underline{\dim}_{B}(F)=\liminf_{\delta\to 0^{+}}\frac{\log N_{\delta}(F)}{-\log\delta},~\text{respectively}.$
Furthermore, if the above two limits exist and are equal, then the box dimension of $F$ is given by
${\dim}_{B}(F)=\lim_{\delta\to 0^{+}}\frac{\log N_{\delta}(F)}{-\log\delta}.$
If $|U_{i}|\subseteq\mathbb{R}^{n}$ represents the diameter of set $U_{i}.$ Then, for $0\leq s\leq n$ the $s$-dimensional Hausdorff measure of a set $F\subseteq\mathbb{R}^{n}$ is defined by
$H^{s}(F)=\lim_{\delta\to\infty}H^{s}_{\delta}(F),$
where $H^{s}_{\delta}(F)=\inf\big{\{}\sum_{i=1}^{\infty}|U_{i}|^{s}:F\subseteq\bigcup_{i=1}^{\infty}U_{i}~\text{and}~|U_{i}|\leq\delta\big{\}}.$ The Hausdorff dimension of $F\subseteq\mathbb{R}^{n}$ is the threshold value at which the Hausdorff measure jumps from infinity to zero, i.e., the Hausdorff dimension of $F$ is the number $0\leq d\leq n$ such that $H^{s}(F)=\infty,$ if $0\leq s<d$ and $H^{d}(F)=0,$ if $d<s\leq n.$
2.1. Code Space
Let $I^{\infty}$ denote the set of all infinite sequences in $\{1,\ldots,N\}$, that is,
$I^{\infty}:=\left\{\{\omega_{k}\}_{k=1}^{\infty}\,\middle|\,\omega_{k}\in\{1,\ldots,N\}\right\}.$
An element of $I^{\infty}$ is written as $\omega:=\omega_{1}\omega_{2}\dots\omega_{k}\omega_{k+1}\dots$, and the $k$-th term of the sequence is denoted by $\omega_{k}$.
For any subset $S\subseteq I^{\infty}$ and a positive integer $n$, define
$S^{n}:=\left\{\omega_{|_{n}}=\omega_{1}\dots\omega_{n}\,\middle|\,\omega=\omega_{1}\omega_{2}\dots\omega_{k}\dots\in S\right\},\quad\text{and}\quad S^{*}:=\bigcup_{n}S^{n}.$
Suppose $(s_{1},s_{2},\ldots,s_{N})$ is a collection of real numbers such that $0<s_{i}<1$ for all $1\leq i\leq N$. A metric $d_{\infty}$ on $I^{\infty}$ is defined by
$d_{\infty}(\omega,\tau):=s_{1}\cdots s_{k},$
if the sequences $\omega=\omega_{1}\omega_{2}\dots\omega_{k}\omega_{k+1}\dots$ and $\tau=\tau_{1}\tau_{2}\dots\tau_{k}\tau_{k+1}\dots$ agree up to the $k$-th term and differ at the $(k+1)$-th term. If the sequences are identical, then $d_{\infty}(\omega,\tau)=0$; and if they differ at the first term, then $d_{\infty}(\omega,\tau)=1$.
Let $\mathcal{I}$ be an iterated function system (IFS) consisting of similarity mappings $\{f_{1},\ldots,f_{N}\}$, and let $A$ be the unique self-similar set associated with $\mathcal{I}$. There is a well-known map $\Theta:I^{\infty}\to A$ defined by
$\Theta(\omega):=\lim_{n\to\infty}f_{\omega_{1}}\circ\cdots\circ f_{\omega_{n}}(x)=\bigcap_{n=1}^{\infty}f_{\omega_{1}}\circ\cdots\circ f_{\omega_{n}}(A),\quad\text{for all}~\omega=\omega_{1}\omega_{2}\dots\in I^{\infty},$
which associates each infinite sequence $\omega$ with a unique point in the attractor $A$.
A compact subset $K\subseteq I^{\infty}$ is said to be closed under left shift if, for every $\omega=\omega_{1}\omega_{2}\dots\in K$, the sequence $\omega^{\prime}:=\omega_{2}\omega_{3}\dots$ also belongs to $K$. Falconer characterizes sub-self-similar sets and establishes a natural correspondence between such sets and subsets of the code space. For more details, see [14].
Proposition 2.1 (see [14], Proposition 2.1).
Let $\mathcal{I}=\{(X,d),f_{1},\ldots,f_{N}\}$ be an IFS. Then a non-empty compact set $F\subseteq X$ is a sub-self-similar set for $\mathcal{I}$ if and only if $F=\Theta(S)$, for some compact set $S\subseteq I^{\infty}$ that is closed under left shift. Here, $\Theta(\omega)=\lim_{n\to\infty}f_{\omega_{1}}\circ\cdots\circ f_{\omega_{n}}(x)$ for any $\omega=\omega_{1}\omega_{2}\dots\in S$ and any $x\in X$.
Remark 2.2.
The convergence $\Theta(\omega)=\lim_{n\to\infty}f_{\omega_{1}\ldots\omega_{n}}(x)$ is independent of the choice of the point $x\in X$.
For any $\omega:=\omega_{1}\dots\omega_{n}\in I^{n}$, we write
$f_{\omega}:=f_{\omega_{1}}\circ f_{\omega_{2}}\dots\circ f_{\omega_{n}},\qquad s_{\omega}:=\prod_{i=1}^{n}s_{\omega_{i}}.$
3. Examples and Construction of Inhomogeneous Sub-Self-Similar Sets
In this section, we present various examples of inhomogeneous sub-self-similar (ISSS) sets and provide one method for their construction.
Example 5.1 (Trivial Cases)
Inhomogeneous self-similar sets and sub-self-similar (SSS) sets are trivial examples of ISSS sets that satisfy Equation (4).
Example 5.2 (Union of ISSS Sets)
Let $F_{C}$ be an ISSS set associated with the IFS $\{f_{1},\ldots,f_{N},C\}$, and let $F_{C^{\prime}}$ be another ISSS set associated with the IFS $\{f^{\prime}_{1},\ldots,f^{\prime}_{N^{\prime}},C^{\prime}\}$. Then the union $F_{C}\cup F_{C^{\prime}}$ is also an ISSS set associated with the IFS $\{f_{1},\ldots,f_{N},f^{\prime}_{1},\ldots,f^{\prime}_{N^{\prime}},C\cup C^{\prime}\}$.
$\displaystyle F_{C}$ $\displaystyle\subseteq\bigcup_{i=1}^{N}f_{i}(F_{C})\cup C,$
$\displaystyle F_{C^{\prime}}$ $\displaystyle\subseteq\bigcup_{i^{\prime}=1}^{N^{\prime}}f^{\prime}_{i^{\prime}}(F_{C^{\prime}})\cup C^{\prime}.$
Thus,
$F_{C}\cup F_{C^{\prime}}\subseteq\bigcup_{i=1}^{N}f_{i}(F_{C}\cup F_{C^{\prime}})\cup\bigcup_{i^{\prime}=1}^{N^{\prime}}f^{\prime}_{i^{\prime}}(F_{C}\cup F_{C^{\prime}})\cup(C\cup C^{\prime}).$
Example 5.3 (Augmenting ISSS Sets)
Let $F_{C}$ be an ISSS set associated with the IFS $\{f_{1},\ldots,f_{N},C\}$. If $K\subseteq F_{C}$ is a compact subset and $f$ is a contracting similarity, then the set $F_{C}\cup f(K)$ is an ISSS set for the IFS $\{f_{1},\ldots,f_{N},f,C\}$. Since $f(K)\subseteq f(F_{C})$, we have:
$F_{C}\cup f(K)\subseteq\bigcup_{i=1}^{N}f_{i}(F_{C}\cup f(K))\cup f(F_{C}\cup f(K))\cup C.$
Example 5.4 (Boundary of an ISSS Set)
Let $F_{C}$ be an ISSS set satisfying Equation (4). Then the topological boundary $\partial F_{C}$ of $F_{C}$ is also an ISSS set. Any neighborhood of $x\in\partial F_{C}$ contains points outside $F_{C}$, implying that $x\in\partial f_{i}(F_{C})\subseteq f_{i}(\partial F_{C})$ or $x\in\partial C\subseteq C$.
Remark 3.1.
Unlike inhomogeneous self-similar sets, ISSS sets are not unique. For example, consider the IFS $\mathcal{I}=\{f_{1}(x)=\frac{x}{2},f_{2}(x)=\frac{x}{2}+\frac{1}{2}\}$ with condensation set $C$. Both the set $\{0,1\}$ and the set of all dyadic numbers in $[0,1]$ are ISSS sets corresponding to $\mathcal{I}$ and $C$.
Remark 3.2.
Let $E$ be a bounded set that is not sub-self-similar with respect to an IFS $\{f_{i}\}_{i=1}^{N}$. Then the set $C=\overline{E\setminus\bigcup_{i=1}^{N}f_{i}(E)}$ ensures that $E=\bigcup_{i=1}^{N}f_{i}(E)\cup C$, making $E$ an ISSS set. Moreover, for any $C^{\prime}\supseteq C$, $E$ remains an ISSS set.
Remark 3.3.
In [31], McClure and Vallin provide examples of non-sub-self-similar sets, such as $E=\{0,1,\frac{1}{2},\frac{1}{3},\ldots\}$, which is not a sub-self-similar set. According to the previous remark, there exists a condensation set $C$ such that $E$ becomes an ISSS set. Furthermore, for any compact set $E$, the minimal condensation set $C=\overline{E\setminus\bigcup_{i=1}^{N}f_{i}(E)}$ ensures that $E$ is an ISSS set.
Remark 3.4.
Inhomogeneous self-similar sets are ISSS sets. Also, if we consider the condensation sets $C$ as the empty set, then ISSS sets reduce to SSS sets, which is a generalization of self-similar sets. In this way, ISSS sets are a generalization of self-similar sets.
4. Construction of Inhomogeneous Sub-Self-Similar Sets
We now present a constructive method for generating ISSS sets from a given sub-self-similar set and a condensation set.
Theorem 4.1.
Let $E$ be a sub-self-similar set associated with the IFS $\mathcal{I}$ such that $\Theta(S)=E$, and let $C\subset X$ be a non-empty compact set. Define
$O_{S}=\bigcup_{{\omega}\in S^{*}}f_{{\omega}}(C).$
Then $E\cup O_{S}$ is an ISSS set corresponding to the IFS $\mathcal{I}$ with condensation set $C$.
Proof.
To prove that $E\cup O_{S}$ is an ISSS set, we first show that it is a compact set. Since $O_{S}$ is bounded, it suffices to prove that $E\cup O_{S}$ is closed. Let $\{x_{n}\}$ be a sequence in $E\cup O_{S}$ converging to some $x\in X$. If $x_{n}\in E$ for infinitely many $n$, then there exists a subsequence $\{x_{n_{r}}\}$ with $x_{n_{r}}\in E$. Since $E$ is closed, $\lim x_{n_{r}}=x\in E\subseteq E\cup O_{S}$. If $x_{n}\in E$ for only finitely many $n$, i.e., $\{x_{n}\}$ is eventually in $O_{S}$, then there exists $n_{0}$ such that $x_{n}\in O_{S}$ for all $n\geq n_{0}$. Hence, for each $n\geq n_{0}$, there exists a string $\omega_{|_{n}}\in S^{*}$ such that $x_{n}\in f_{\omega_{|_{n}}}(C)$. Let us denote the length of any string $\omega:=\omega_{1}\omega_{2}\dots\omega_{k}\in S^{*}$ by $|\omega|=k.$
Case 1: $\liminf|\omega_{|_{n}}|=k<\infty$. Then there exists a subsequence $\{\omega_{{|_{n}}_{j}}\}$ of $\{\omega_{|_{n}}\}$ such that $\omega_{{|_{n}}_{j}}\in S^{k}$ for all $j$, where $S^{k}$ is a finite set. Thus, there exists a string $t\in\{\omega_{{|_{n}}_{j}}\mid j\in\mathbb{N}\}$ and a further subsequence $\{\omega_{{|_{n}}_{j_{p}}}\}$ such that $\omega_{{|_{n}}_{j_{p}}}=t$ for all $p$. Then $x_{n_{j_{p}}}\in f_{t}(C)$, and since $f_{t}(C)$ is closed, $\lim x_{n_{j_{p}}}=x\in f_{t}(C)\subseteq O_{S}$.
Case 2: $\liminf|\omega_{|_{n}}|=\infty$. Then
$\operatorname{diam}(f_{\omega_{|_{n}}}(C)\cup f_{\omega_{|_{n}}}(E))$ $=\rho_{\omega_{|_{n}}}\operatorname{diam}(C\cup E)$
$\leq\rho_{\max}^{|\omega_{|_{n}}|}\operatorname{diam}(C\cup E)\to 0\quad\text{as }n\to\infty,$
since $\rho_{\max}=\max\{\rho_{i}\}<1$.
Now, for each $n$, choose $y_{n}\in f_{\omega_{|_{n}}}(E)$. For positive integers $m$ and $n$,
$|y_{n}-y_{m}|$ $\leq|y_{n}-x_{n}|+|x_{n}-x_{m}|+|x_{m}-y_{m}|$
$\leq\operatorname{diam}(f_{\omega_{|_{n}}}(C)\cup f_{\omega_{|_{n}}}(E))+|x_{n}-x_{m}|+\operatorname{diam}(f_{\omega_{|_{m}}}(C)\cup f_{\omega_{|_{m}}}(E)).$
Hence, $\{y_{n}\}$ is a Cauchy sequence and converges to some $y\in X$. Then,
$|x-y|$ $\leq|x-x_{n}|+|x_{n}-y_{n}|+|y_{n}-y|$
$\leq|x-x_{n}|+\operatorname{diam}(f_{\omega_{|_{n}}}(C)\cup f_{\omega_{|_{n}}}(E))+|y_{n}-y|\to 0.$
Thus, $x=y$.
Since $y_{n}\in f_{\omega_{|_{n}}}(E)$ for each $n$, there exists $y^{\prime}\in E$ such that $y_{n}=f_{\omega_{1}\ldots\omega_{n}}(y^{\prime})$. Let $\epsilon>0$. Choose $n_{1}\in\mathbb{N}$ such that for all $n\geq n_{1}$,
$|y_{n}-y|=|f_{\omega_{1}\ldots\omega_{n}}(y^{\prime})-y|<\epsilon/2.$
Since $\Theta(S)=E$, we have $\lim_{n\to\infty}f_{\omega_{1}\ldots\omega_{n}}(y^{\prime})=z_{n_{\epsilon}}\in E$. Choose $n_{2}\in\mathbb{N}$ such that for all $n\geq n_{2}$,
$|f_{\omega_{1}\ldots\omega_{n}}(y^{\prime})-z_{n_{\epsilon}}|<\epsilon/2.$
Let $n^{\prime}=\max\{n_{1},n_{2}\}$. Then,
$|y-z_{n_{\epsilon}}|$ $\leq|f_{\omega_{1}\ldots\omega_{n}}(y^{\prime})-y|+|f_{\omega_{1}\ldots\omega_{n}}(y^{\prime})-z_{n_{\epsilon}}|$
$<\epsilon/2+\epsilon/2=\epsilon.$
Since $E$ is closed, $x=y=\lim z_{n_{\epsilon}}\in E\subseteq E\cup O_{S}$. Thus, $E\cup O_{S}$ is closed, and hence compact.
It remains to show that $E\cup O_{S}$ satisfies Equation (4).
Clearly,
$E\cup O_{S}$ $\subseteq\bigcup_{i=1}^{N}f_{i}(E)\cup O_{S}$
$\subseteq\bigcup_{i=1}^{N}f_{i}(E\cup O_{S})\cup O_{S}\cup C.$
Now we show that $O_{S}\subseteq\bigcup_{i=1}^{N}f_{i}(O_{S})$. Let $x\in O_{S}$. Then there exists $\omega_{|_{n}}\in S^{*}$ and $c\in C$ such that
$x=f_{\omega_{|_{n}}}(c)=f_{\omega_{1}\ldots\omega_{n}}(c).$
Let $x^{\prime}=f_{\omega_{2}\ldots\omega_{n}}(c)\in O_{S}$. Then,
$x=f_{\omega_{1}}(x^{\prime})\in\bigcup_{i=1}^{N}f_{i}(O_{S}).$
Thus,
$E\cup O_{S}\subseteq\bigcup_{i=1}^{N}f_{i}(E\cup O_{S})\cup C.$
This completes the proof. ∎
Theorem 4.2.
Let $E$ be a sub-self-similar set with $\Theta(S)=E$, and let
$O_{S}=\bigcup_{\omega\in S^{*}}f_{\omega}(C).$
Then,
$\overline{O}_{S}=E\cup O_{S}.$
Proof.
By the construction, $\overline{O}_{S}$ is the smallest closed set containing $O_{S}$. Since $E\cup O_{S}$ is closed and $O_{S}\subseteq E\cup O_{S}$, it follows that $\overline{O}_{S}\subseteq E\cup O_{S}$.
To prove the reverse inclusion, let $x\in E\cup O_{S}$ be any element. If $x\in O_{S}$, then nothing to prove, if $x\in E$, then by 2.1, there exists a sequence $\omega=(\omega_{1},\omega_{2},\ldots)\in S$ such that $x=\Theta(\omega)$. We will construct a sequence $\{x_{n}\}$ in $O_{S}$ which converges to $x\in E.$ Fix $c\in C$ and define $x_{n}=f_{\omega_{1}\cdots\omega_{n}}(c)\in O_{S}$. Then $x_{n}\to x\in E$, so $x\in\overline{O}_{S}$. Thus, $E\cup O_{S}\subseteq\overline{O}_{S}$. ∎
5. Dimension of inhomogeneous sub-self-similar sets
In this section, we aim to estimate the Hausdorff and box dimensions for ISSS sets (arrived from a given SSS set $E$ and a condensation set $C$), which generalize to the inhomogeneous self-similar sets.
5.1. Hausdorff dimension
Theorem 5.1.
(see,  [14, Theorem  3.5]) Let $E$ be a SSS set for a family of contracting similarities $f_{1},\ldots,f_{N}$. Let $s$ be the unique non-negative number such that $\tau(s)\equiv\lim_{k\to\infty}\bigg{(}\sum_{i\in S_{k}}\rho_{i}^{s}\bigg{)}^{1/k}=1,$ then
$\dim_{H}E\leq\underline{\dim}_{B}E\leq\overline{\dim}_{B}E\leq s.$
Furthermore, if it satisfies the open set condition, then
$\dim_{H}E=\underline{\dim}_{B}E=\overline{\dim}_{B}E=s.$
Corollary 5.2.
Let $E\cup O_{S}$ be an ISSS set with SSS set $E$ and condensation set $C.$ Then $E\cup O_{S}=\max\{s,\dim_{H}C\}$.
Proof.
Since the Hausdorff dimension is countably stable, the Hausdorff dimension of $O_{S}$ equals the Hausdorff dimension of $C$. Additionally, the Hausdorff dimension of the union of $E\cup O_{S}=\max\{s,\dim_{H}C\}$. ∎
5.2. Upper box dimension
In order to estimate the box dimension of these ISSS sets, we will follow Fraser’s $\delta$-covering technique. For more details, one can see [15].
Lemma 5.3.
If $t>s,$ then there exists a constant $m_{t}$ (depends only on $t$) such that
$\sum_{\omega\in S^{*}}\rho_{\omega}^{t}=m_{t}<\infty.$
Proof.
The function $\tau(s)=\lim_{k\to\infty}\bigg{(}\sum_{\omega\in S^{k}}\rho_{\omega}^{s}\bigg{)}^{1/k}$ defined in Theorem (5.1) is strictly decreasing continuous function with $0\leq\tau(s)<\infty$ and $\sum_{\omega\in S^{k}}\rho_{\omega}^{s}\geq\tau(s)^{k}$ for all $k=1,2,\ldots$
Also observe that $\sum_{\omega\in S^{*}}\rho_{\omega}^{t}=\sum_{k=1}^{\infty}\sum_{\omega\in S^{k}}\rho_{\omega}^{t}$ converges if $\tau(s)<1$ and diverges if $\tau(s)>1.$ Hence for $t>s$$\sum_{\omega\in S^{*}}\rho_{\omega}^{t}=\sum_{k=1}^{\infty}\sum_{\omega\in S^{k}}\rho_{\omega}^{t}=m_{t}<\infty,$ completing the assertion. ∎
For $\delta\in(0,1],$ defining $\delta$-stopping, for $S$ as $S(\delta)=\{\omega\in S^{*}:\rho_{\omega}<\delta\leq\rho_{\omega_{-}}\}.$ Where $\omega_{-}$ is the prefix of $\omega$ with $|\omega_{-}|=n-1,$ if $|\omega|=n.$ In the following lemma, we estimate a bound for $|S(\delta)|.$
Lemma 5.4.
If $t>s,$ then
$|S(\delta)|\leq m_{t}\rho_{\min}^{-t}\delta^{-t}.$
For all $\delta\in(0,1].$ Where $\rho_{\min}=\min_{i=1}^{N}\{\rho_{i}\}$
Proof.
By the above lemma, we have $m_{t}=\sum_{\omega\in S^{*}}\rho_{\omega}^{t}\geq\sum_{\omega\in S(\delta)}\rho_{\omega}^{t}.$
Also for $\omega\in S(\delta),\rho_{\omega}\geq\rho_{\omega_{-}}\rho_{\min}\geq\delta\rho_{\min}.$ Using this, we get
$m_{t}\geq\sum_{\omega\in S(\delta)}\delta^{t}\rho_{\min}^{t}=|S(\delta)|\delta^{t}\rho_{\min}^{t}.$
Which completes the proof. ∎
Lemma 5.5.
For $t>s$ and $\delta\in(0,1)$ we have
$|\{\omega\in S^{*}:\delta\leq\rho_{\omega}\}|\leq\frac{\log\delta}{\log\rho_{\max}}\delta^{-t}m_{t}^{l}.$
where $l$ is largest integer such that $\delta\leq\rho_{\max}^{l}.$
Proof.
Let $\delta\in(0,1)$ and $\omega\in S^{*}$ such that $\delta\leq\rho_{\omega}.$ It follows that $\delta\leq\rho_{\max}^{|\omega|}$ and hence we will get $|\omega|\leq\frac{\log\delta}{\log\rho_{\max}}.$ Also, we have
${m_{t}}\frac{\log\delta}{\log\rho_{\max}}$ $\geq\sum_{N\in\mathbb{N},N<\frac{\log\delta}{\log\rho_{\max}}}m_{t}$
$\geq\sum_{N\in\mathbb{N},N<\frac{\log\delta}{\log\rho_{\max}}}\sum_{\omega\in S^{N}}\rho_{\omega}^{t}$
$\geq\sum_{N\in\mathbb{N},N<\frac{\log\delta}{\log\rho_{\max}}}\sum_{\omega\in S^{N},\delta\leq\rho_{\omega}}\delta^{t}$
$=|\{\omega\in S^{*}:\delta\leq\rho_{\omega}\}|\delta^{t}.$
Lemma 5.6.
For $\delta\in(0,1],$ we have
$\bigcup_{\omega\in S^{*},\rho_{\omega}<\delta}f_{\omega}(C)\subseteq\bigcup_{\omega\in S(\delta)}f_{\omega}(X).$
Proof.
For any $x\in\bigcup_{\omega\in S^{*},\rho_{\omega}<\delta}f_{\omega}(C),$ there is some string $\omega=\omega_{1}\ldots\omega_{n}\in S^{*}$ and $c\in C$ such that $S_{\omega}(c)=x$ with $\rho_{\omega}<\delta.$ If we assume that the contraction ratio for empty string is $1,$ then clearly there exists some prefix $\omega^{\prime}=\omega_{1}\ldots\omega_{m}$ of $\omega$ such that $\omega^{\prime}\in S(\delta)$ and $x=f_{\omega^{\prime}}(f_{\omega_{m+1}\ldots\omega_{n}}(c))\in f_{\omega^{\prime}}(X).$ Completing the proof of Lemma. ∎
Theorem 5.7.
Let $E$ be an SSS set with contracting similarities $\{f_{1},\ldots,f_{N}\}$ and $E\cup O_{S}$ be the corresponding ISSS sets with compact set $C.$ Then
$\max\{\overline{\dim}_{B}E,\overline{\dim}_{B}C\}\leq\overline{\dim}_{B}E_{C}\leq\max\{s,\overline{\dim}_{B}C\}.$
Proof.
Since the upper box dimension is finitely stable and monotonic, we have
$\max\{\overline{\dim}_{B}E,\overline{\dim}_{B}C\}\leq\overline{\dim}_{B}E_{C}\leq\max\{\overline{\dim}_{B}E,\overline{\dim}_{B}O_{S}\}.$
By Theorem (5.1) $\overline{\dim}_{B}E\leq s,$ it is sufficient to show that $\overline{\dim}_{B}O_{S}\leq\max\{s,\overline{\dim}_{B}C\}.$ Let $\max\{s,\overline{\dim}_{B}C\}<t,$ then by the definition of the box-dimension, there exists some constant $c_{t}$ such that
$N_{\delta}(C)\leq c_{t}\delta^{-t},$
for all $\delta\in(0,1].$ Consider
$N_{\delta}(O_{S})$ $=N_{\delta}\bigg{(}C\cup\bigcup_{\omega\in S^{*}}f_{\omega}(C)\bigg{)}$
$\leq N_{\delta}(C)+N_{\delta}\bigg{(}\bigcup_{\omega\in S^{*},\rho_{\omega}\geq\delta}f_{\omega}(C)\bigg{)}+N_{\delta}\bigg{(}\bigcup_{\omega\in S^{*},\rho_{\omega}<\delta}f_{\omega}(C)\bigg{)}$
$\leq N_{\delta}(C)+N_{\delta}\bigg{(}\bigcup_{\omega\in S^{*},\rho_{\omega}\geq\delta}f_{\omega}(C)\bigg{)}+N_{\delta}\bigg{(}\bigcup_{\omega\in S(\delta)}f_{\omega}(X)\bigg{)}$
$\leq N_{\delta}(C)+\sum_{\omega\in S^{*},\rho_{\omega}\geq\delta}N_{\delta}(f_{\omega}(C))+\sum_{\omega\in S(\delta)}N_{\delta}(f_{\omega}(X))$
$\leq N_{\delta}(C)+\sum_{\omega\in S^{*},\rho_{\omega}\geq\delta}N_{{\delta}/{\rho_{\omega}}}(C)+\sum_{\omega\in S(\delta)}N_{{\delta}/{\rho_{\omega}}}(X)$
$\leq c_{t}\delta^{-t}+\sum_{\omega\in S^{*},\rho_{\omega}\geq\delta}c_{t}({\delta}/{\rho_{\omega}})^{-t}+\sum_{\omega\in S(\delta)}N_{1}(X)$
$\leq c_{t}\delta^{-t}+c_{t}\delta^{-t}\sum_{\omega\in S^{*}}{\rho_{\omega}}^{t}+|S(\delta)|N_{1}(X)$
$\leq c_{t}\delta^{-t}+c_{t}\delta^{-t}m_{t}+m_{t}{\rho}^{-t}_{\min}\delta^{-t}N_{1}(X)$
$\leq(c_{t}+c_{t}m_{t}+m_{t}{\rho}^{-t}_{\min}N_{1}(X))\delta^{-t}.$
The third line comes from Lemma (5.6). The sixth line comes as $X$ is a compact space and $\delta/\rho_{\omega}>1,$ the last line arrived by Lemma (5.4). This completes the proof. ∎
5.3. Lower box dimension
In this subsection, our goal is to estimate the upper bound for the lower box dimension of the ISSS sets. We observe that if $\overline{\dim}_{B}C=\underline{\dim}_{B}C$, then the box dimension of ISSS sets is equal to the maximum of the dimension of the corresponding homogeneous SSS set and the box dimension of the condensation set. Since the lower box dimension follows the monotonicity property, we have
$\max\left\{\underline{\dim}_{B}E,\underline{\dim}_{B}C\right\}\leq\underline{\dim}_{B}(E\cup O_{S})\leq\max\left\{s,\overline{\dim}_{B}C\right\}.$
Given the stability of the lower box dimension under closure, we have
$\underline{\dim}_{B}(E\cup O_{S})=\underline{\dim}_{B}\overline{O}_{S}=\underline{\dim}_{B}O_{S}.$
We will use the concept of the covering regularity index of the condensation set $C$ from [15] to estimate an upper bound for the lower box dimension of the set $O_{S}$.
For $t\geq 0$ and $\delta\in(0,1]$, the $(t,\delta)$ covering regularity exponent of $C$ is defined as
$P_{t,\delta}(C)=\sup\left\{p\in[0,1]:N_{\delta^{p}}(C)\geq\delta^{-pt}\right\}.$
The $t$-covering regularity exponent of $C$ is defined as
$P_{t}(C)=\liminf_{\delta\to 0}P_{t,\delta}(C).$
Lemma 5.8.
(see  [15, Lemma   2.6]) For any $C\in\mathbb{R}^{n}$, we have:
  1. (1)
    For all $t,\delta>0$, $P_{t,\delta}(C)$, $P_{t}(C)\in[0,1]$.
  2. (2)
    $P_{t}(C)$ is decreasing in $t$, and if $t<\underline{\dim}_{B}C$, then $P_{t}(C)=1$, and if $t>\overline{\dim}_{B}C$, then $P_{t}(C)=0$.
  3. (3)
    For all $t>\underline{\dim}_{B}C$, we have $P_{t}(C)\leq\underline{\dim}_{B}C<t$.
Lemma 5.9.
(see,[15, Lemma  3.6]) For $t\geq 0$, if $P_{t}(C)<1$, then for all $\epsilon\in(0,1-P_{t}(C))$, there exists $\delta\in(0,\epsilon)$ such that:
$P_{t}(C)-\epsilon<P_{t,\delta}(C)<P_{t}(C)+\epsilon,$
and for all $\delta_{0}\in[\delta,\delta^{P_{t}(C)}]$, we have:
$N_{\delta_{0}}(C)\leq\delta_{0}^{-t}.$
By Lemma (5.4), for $t>s$, the cardinality of the $\delta$-stopping set $S$ is given by:
$|S(\delta)|\leq m_{t}\rho_{\min}^{-t}\delta^{-t}.$
Using the above two lemmas, we can estimate the lower box dimension.
Theorem 5.10.
For all $t>\max\{s,\underline{\dim}_{B}C\}$, we have:
$\underline{\dim}_{B}(E\cup O_{S})\leq\max\{s,\underline{\dim}_{B}C\}.$
Proof.
Let $t>\max\{s,\underline{\dim}_{B}C\}$. Then, by Lemma (5.8), we have $P_{t}(C)\leq\underline{\dim}_{B}C<t$. Thus, by Lemma (5.9), for all $\epsilon\in(0,1-P_{t}(C))$, there exists $\delta\in(0,\epsilon)$ such that:
$P_{t}(C)-\epsilon<P_{t,\delta}(C)<P_{t}(C)+\epsilon.$
Moreover, for all $\delta_{0}\in[\delta,\delta^{P_{t}(C)}]$, we have:
$N_{\delta_{0}}\leq\delta_{0}^{-t}.$
For a fixed $\epsilon\in(0,1-P_{t}(C))$ and choosing $\delta\in(0,\epsilon)$, we have:
$N_{\delta}(O_{S})$ $=N_{\delta}\left(C\cup\bigcup_{\omega\in S^{*}}f_{\omega}(C)\right)$
$\leq\sum_{\omega\in S^{*},\delta^{1-P_{t,\delta}(C)-\epsilon}\leq\rho_{\omega}<1}N_{\delta}f_{\omega}(C)+\sum_{\omega\in S^{*},\delta\leq\rho_{\omega}<\delta^{1-P_{t,\delta}(C)-\epsilon}}N_{\delta}f_{\omega}(C)$
$\quad\hskip 128.0374pt+N_{\delta}\left(\bigcup_{\omega\in S^{*},\rho_{\omega}<\delta}f_{\omega}(C)\right)+N_{\delta}(C)$
$\leq\sum_{\omega\in S^{*},\delta^{1-P_{t,\delta}(C)-\epsilon}\leq\rho_{\omega}<1}N_{\delta/\rho_{\omega}}(C)+\sum_{\omega\in S^{*},\delta\leq\rho_{\omega}<\delta^{1-P_{t,\delta}(C)-\epsilon}}N_{\delta/\rho_{\omega}}(C)$
$\quad\hskip 125.19194pt+N_{\delta}\left(\bigcup_{\omega\in S(\delta),\rho_{\omega}<\delta}f_{\omega}(X)\right)+N_{\delta}(C)$
$\leq\sum_{\omega\in S^{*},\delta^{1-P_{t,\delta}(C)-\epsilon}\leq\rho_{\omega}<1}(\delta/\rho_{\omega})^{-t}+\sum_{\omega\in S^{*},\delta\leq\rho_{\omega}<\delta^{1-P_{t,\delta}(C)-\epsilon}}N_{\delta^{P_{t,\delta}(C)+\epsilon}}(C)$
$\quad\hskip 128.0374pt+\sum_{\omega\in S(\delta)}N_{\delta/\rho_{\omega}}(X)+\delta^{-t}$
$\leq\sum_{\omega\in S^{*},\delta^{1-P_{t,\delta}(C)-\epsilon}\leq\rho_{\omega}<1}(\delta/\rho_{\omega})^{-t}+\sum_{\omega\in S^{*},\delta\leq\rho_{\omega}<\delta^{1-P_{t,\delta}(C)-\epsilon}}N_{\delta/\rho_{\omega}}(C)$
$\quad\hskip 128.0374pt+N_{1}(X)|S(\delta)|+\delta^{-t}$
$\leq\sum_{\omega\in S^{*}}(\delta/\rho_{\omega})^{-t}+|\{\omega\in S^{*}:\delta\leq\rho_{\omega}\}|\delta^{-(P_{t,\delta}(C)+\epsilon)t}$
$\quad\hskip 128.0374pt+N_{1}(X)m_{t}\rho_{\min}^{-t}\delta^{-t}+\delta^{-t}$
$\leq\delta^{-t}\sum_{\omega\in S^{*}}\rho_{\omega}^{t}+|\{\omega\in S^{*}:\delta\leq\rho_{\omega}\}|\delta^{-(P_{t,\delta}(C)+\epsilon)t}$
$\quad\hskip 128.0374pt+N_{1}(X)m_{t}\rho_{\min}^{-t}\delta^{-t}+\delta^{-t}$
$\leq\left(m_{t}+N_{1}(X)m_{t}\rho_{\min}^{-t}+1+m_{t}\frac{\log\delta}{\log\rho_{\max}}\delta^{-(P_{t,\delta}(C)+\epsilon)t}\right)\delta^{-t}$
$\leq\left(m_{t}+N_{1}(X)m_{t}\rho_{\min}^{-t}+1+m_{t}\frac{\log\delta}{\log\rho_{\max}}\delta^{-(P_{t}(C)+2\epsilon)t}\right)\delta^{-t}.$
Letting $\epsilon\to 0$, we get $\underline{\dim}_{B}O_{S}\leq t$. This concludes $\underline{\dim}_{B}(E\cup O_{S})\leq\max\{s,\underline{\dim}_{B}C\}$.
Remark 5.11.
It should be noted that if we take $S=I^{\infty}$ (the code space), then the given sub-self-similar set reduces to the self-similar set associated with the underlying IFS. In this case, the above dimension estimation yields the Hausdorff and box dimensions of the inhomogeneous self-similar set associated with the condensation set $C$.
6. Continuity of the Hausdorff dimension of ISSS
In this section, we observe that for each ISSS set of the form $F:=E\cup O_{S}$, where $E$ is an SSS set and $C$ is the condensation set, there exists a sequence of inhomogeneous IFSs $\{I_{k}\}$ with corresponding ISSS sets $\{F_{k}\}$ such that $F=\bigcup_{k\in\mathbb{N}}F_{k}$. Furthermore, we discuss the continuity of the Hausdorff dimension of ISSS sets.
Theorem 6.1.
Let $F=E\cup O_{S}$ be the ISSS set corresponding to the IFS $\{f_{i}\}_{i=1}^{N}$ with condensation set $C$, which satisfies the open set condition. Then, there exists a sequence of IFSs $\{I_{k}\}_{k=1}^{\infty}$ such that
$F=\bigcup_{k=1}^{\infty}F_{k},$
where $F_{k}=E_{k}\cup O_{S}$ are the ISSS sets corresponding to the IFSs $I_{k}$, each with condensation set $C$.
Proof.
Let $I_{k}=\{f_{\omega}:\omega\in S^{k}\}$. For any $x\in F=E\cup O_{S}$, we consider two cases:
Case 1: If $x\in E$, then there exists a string $\omega=\omega_{1}\omega_{2}\ldots\in S$ such that
$x=\lim_{n\to\infty}f_{\omega_{1}}\circ f_{\omega_{2}}\circ\ldots\circ f_{\omega_{n}}(z).$
For each $k\in\mathbb{N}$, define $\omega_{j}|_{k}^{k}:=\omega_{jk+1}\omega_{jk+2}\ldots\omega_{jk+k}$ for $j=0,1,2,\ldots$. Then $\omega_{j}|_{k}^{k}\in S^{k}$, and we have
$x=\lim_{j\to\infty}f_{\omega_{0}|_{k}^{k}}\circ f_{\omega_{1}|_{k}^{k}}\circ\ldots\circ f_{\omega_{j}|_{k}^{k}}(z),$
which implies that $x\in\bigcup_{k=1}^{\infty}F_{k}$.
Case 2: If $x\in O_{S}$, then there exists some $\omega=\omega_{1}\omega_{2}\ldots\omega_{n}\in S^{*}$ and $c\in C$ such that
$f_{\omega_{1}}\circ\ldots\circ f_{\omega_{n}}(c)=x.$
Clearly, $\omega_{1}\ldots\omega_{n}\in S^{n}$, so $x\in\bigcup_{k=1}^{\infty}F_{k}\cup O_{S}$.
Conversely, suppose $x\in\bigcup_{k=1}^{\infty}F_{k}\cup O_{S}$. If $x\in O_{S}$, then the claim holds trivially. Now suppose $x\in F_{k}$ for some $k$. Then there exists a sequence $\omega_{0}|_{k}^{k},\omega_{1}|_{k}^{k},\ldots,\omega_{j}|_{k}^{k}\in S^{k}$ such that
$x=\lim_{j\to\infty}f_{\omega_{0}|_{k}^{k}}\circ f_{\omega_{1}|_{k}^{k}}\circ\ldots\circ f_{\omega_{j}|_{k}^{k}}(z).$
Using the fact that $\Theta(S)=E$ and that the concatenation $\omega=\omega_{0}|_{k}^{k}\omega_{1}|_{k}^{k}\ldots\omega_{j}|_{k}^{k}\in S$, we conclude that $x\in F$.
Thus, we obtain $F=\bigcup_{k=1}^{\infty}F_{k}$. ∎
Lemma 6.2.
[27] Let $S\subseteq l^{\infty}$ be any shift-invariant set in code space $I^{\infty}.$ Then
$\dim_{H}S=\dim_{B}S=\lim_{k\to\infty}s_{k},$
where the relation
$\sum_{\omega\in S^{k}}(\rho_{\omega})^{s_{k}}=1$
determines $s_{k}.$
Corollary 6.3.
Let $F=E\cup O_{S}$ be the ISSS set corresponding to the IFS $\{f_{i}\}{i=1}^{N}.$ Then there exists a sequence of IFSs $\{I_{k}\}$ with ISSS sets $\{F_{k}\}$ and under inhomogeneous open set condition (IOSC) $\dim_{H}E=\lim_{k\to\infty}\dim_{H}E_{k}=\lim_{k\to\infty}=\lim_{k\to\infty}s_{k},$ where $s_{k}$ is the Hausdorff dimension of $F_{k}$ defined in Lemma 6.2.
Proof.
By Theorem 6.1, there exists a sequence of IFSs $\{I_{k}\}$ with ISSS sets $F_{k}$ such that $F=\bigcup_{k=1}^{\infty}F_{k}.$ It is easy to observe that if $\{f_{i}\}_{i=1}^{N}$ satisfies the IOSC, then each $I_{k}$ also satisfies the IOSC. If $s_{k}$ represents the Hausdorff dimension of $E_{k}$, which satisfies
$\lim_{k\to\infty}\left(\sum_{\omega\in S^{k}}\rho_{\omega}^{s_{k}}\right)^{1/k}=1,$
then the Hausdorff dimension of $F_{k}$ is $\max\{s_{k},\dim_{H}(C)\}$.
By Lemma (6.2), we have
$\dim_{H}E=\lim_{k\to\infty}\dim_{H}E_{k}=\lim_{k\to\infty}s_{k},$
and $\dim_{H}F_{k}\to\dim_{H}F.$ This completes the assertion. ∎
7. Self-similar structure on the product of IFS
In this section, we define the product of inhomogeneous iterated function systems (IFSs) and examine the associated separation conditions. We conclude by demonstrating that the product of the inhomogeneous invariant measures corresponding to the individual IFSs constitutes an inhomogeneous type invariant measure for the product IFS. Let $(X,d_{1})$ and $(Y,d_{2})$ be two complete metric spaces; then it is well known that the product space $(X\times Y,d)$ is a complete metric space with respect to the metric $d$ given by
$d((x,y),(z,w))=\max\{d_{1}(x,z),d_{2}(y,w)\}.$
Now, for two IFSs $\mathcal{I}_{1}=\{(X,d_{1});f_{1},f_{2},\ldots,f_{N}\}$ and $\mathcal{I}_{2}=\{(Y,d_{2});g_{1},g_{2},\ldots,g_{M}\}$ we define product IFS $\mathcal{P}=\{X\times Y;\Phi_{(ij)},(ij)\in\{1,2,\ldots,N\}\times\{1,2,\ldots,M\}\}$, where $\Phi_{(ij)}(x,y)=(f_{i}(x),g_{j}(y)).$ It is easy to observe that $\Phi_{ij}$ is contraction for all $ij\in\{1,2,\ldots,N\}\times\{1,2,\ldots,M\}.$ For any $(x,y),(z,w)\in X\times Y,$ we have
$d(\Phi_{ij}(x,y),\Phi_{ij}(z,w))$ $=d((f_{i}(x),g_{j}(y),(f_{i}(z),g_{j}(w)))$
$=\max(d_{1}(f_{i}(x),f_{i}(z)),d_{2}(g_{j}(y),g_{j}(w)))$
$\leq\{c^{1}_{i}d_{1}(x,z),c^{2}_{j}d_{2}(z,w)\}$
$\leq c\max\{d_{1}(x,z),d_{2}(z,w)\}$
$=c~d((x,y),(z,w)),$
where $c=\max\{c^{1}_{i},c^{2}_{j}\}.$ Let $A_{\mathcal{I}_{1}}$ and $A_{\mathcal{I}_{2}}$ be the attractors corresponding to the IFSs ${\mathcal{I}_{1}}$ and ${\mathcal{I}_{2}},$ respectively then It is easy to verify that the attractor corresponding to the product IFS $\mathcal{P}$ is $A_{\mathcal{I}_{1}}\times A_{\mathcal{I}_{2}}.$ We claim that $\bigcup_{i,j}f_{i}(A_{\mathcal{I}_{1}})\times g_{j}(A_{\mathcal{I}_{2}})=\bigcup_{i}f_{i}(A)\times\bigcup_{j}g_{j}(A_{\mathcal{I}_{2}}).$
Let $(x,y)\in\bigcup_{i,j}f_{i}(A_{\mathcal{I}_{1}})\times g_{j}(A_{\mathcal{I}_{2}})$
$\Longleftrightarrow(x,y)\in f_{i}(A_{\mathcal{I}_{1}})\times g_{j}(A_{\mathcal{I}_{2}})~\text{for some}~(ij)\in\{1,2,\ldots,N\}\times\{1,2,\ldots,M\}$
$\Longleftrightarrow(x,y)\in f_{i}(A_{\mathcal{I}_{1}})\times g_{j}(A_{\mathcal{I}_{2}})$
$\Longleftrightarrow x\in f_{i}(A_{\mathcal{I}_{1}})~\text{and}~y\in g_{j}(A_{\mathcal{I}_{2}})$
$\Longleftrightarrow x\in\bigcup_{i}f_{i}(A_{\mathcal{I}_{1}})~\text{and}~y\in\bigcup_{j}g_{j}(A_{\mathcal{I}_{2}})$
$\Longleftrightarrow(x,y)\in\bigcup_{i}f_{i}(A_{\mathcal{I}_{1}})\times\bigcup_{j}g_{j}(A_{\mathcal{I}_{2}}).$
Now,
$A_{\mathcal{I}_{1}}\times A_{\mathcal{I}_{2}}$ $=\bigcup_{i}f_{i}(A_{\mathcal{I}_{1}})\times\bigcup_{i}g_{j}(A_{\mathcal{I}_{1}})$
$=\bigcup_{i,j}f_{i}(A_{\mathcal{I}_{1}})\times g_{j}(A_{\mathcal{I}_{2}}).$
In the similar fashion, we define the product of two inhomogeneous IFS. Let $\mathcal{I}_{C_{1}}=\{(X,d_{1});f_{1},f_{2},\ldots,f_{N},C_{1}\}$ and $\mathcal{I}_{C_{2}}=\{(Y,d_{2});g_{1},g_{2},\ldots,g_{M},C_{2}\}$ be two inhomogeneous IFSs with condensation set $C_{1}\subset X$ and $C_{2}\subset Y$ respectively, then we define inhomogeneous product IFS $\mathcal{P}_{C}=\{X\times Y;\Phi_{(ij)},(C_{1},C_{2}),(ij)\in\{1,2,\ldots,N\}\times\{1,2,\ldots,M\}\}.$ If $A_{C_{1}}$ and $A_{C_{2}}$ are the inhomogeneous attractors corresponding to the IFSs $\mathcal{I}_{C_{1}}$ and $\mathcal{I}_{C_{2}},$ respectively. Then $A_{C_{1}}\times A_{C_{2}}$ is a super-self-similar set corresponding to the product IFS $\mathcal{P}_{C},$ i.e., $A_{C_{1}}\times A_{C_{2}}\supseteq\bigcup_{i,j}\Phi_{ij}(A_{C_{1}}\times A_{C_{2}})\cup C_{1}\times C_{2}.$
Theorem 7.1.
Let $A_{C_{1}}$ and $A_{C_{2}}$ be the inhomogeneous attractors corresponding to the inhomogeneous IFSs $\mathcal{I}_{C_{1}}=\{(X,d_{1});f_{1},f_{2},\ldots,f_{N},C_{1}\}$ and $\mathcal{I}_{C_{2}}=\{(Y,d_{2});g_{1},g_{2},\ldots,g_{M},C_{2}\},$ respectively. Let $\mathcal{P}_{C}=\{X\times Y;\Phi_{(ij)},(C_{1},C_{2}),(ij)\in\{1,2,\ldots,N\}\times\{1,2,\ldots,M\}\}$ be the product IFS, then $A_{C_{1}}\times A_{C_{2}}$ be a super-self-similar set.
Proof.
Since $A_{C_{1}}$ and $A_{C_{2}}$ are the inhomogeneous attractors. We have
$A_{C_{1}}=\bigcup_{i=1}^{N}f_{i}(A_{c_{1}})\cup C_{1},~\text{and}~A_{C_{2}}=\bigcup_{j=1}^{M}g_{j}(A_{C_{2}})\cup C_{2}.$
Also,
$A_{C_{1}}\times A_{C_{2}}$ $=\bigg{(}\bigcup_{i}f_{i}(A_{C_{1}})\cup C_{1}\bigg{)}\times\bigg{(}\bigcup_{j}g_{j}(A_{C_{2}})\cup C_{2}\bigg{)}$
$=\bigg{(}\bigcup_{i}f_{i}(A_{C_{1}})\times\bigcup_{j}g_{j}(A_{C_{2}})\bigg{)}\cup\bigg{(}\bigcup_{i}f_{i}(A_{C_{2}})\times C_{2}\bigg{)}\cup$
$\hskip 113.81102pt\bigg{(}C_{1}\times\bigcup_{j}g_{j}(A_{C_{2}})\bigg{)}\cup(C_{1}\times C_{2})$
$\supseteq\bigcup_{i,j}f_{i}(A_{C_{1}})\times g_{j}(A_{C_{2}})\cup(C_{1}\times C_{2}).$
In the next theorem, we see that the inhomogeneous separation conditions on $\mathcal{I}_{C_{1}}$ and $\mathcal{I}_{C_{2}}$ can be transferred to the inhomogeneous product IFS $\mathcal{P}_{C}$.
Theorem 7.2.
Let $\mathcal{I}_{1}=\{(X,d_{1});f_{1},f_{2},\ldots,f_{N},C_{1}\}$ and $\mathcal{I}_{2}=\{(Y,d_{2});g_{1},g_{2},\ldots,g_{M},C_{2}\}$ be two IFSs. Then the following hold:
  1. (1)
    The IFS $\mathcal{P}_{C}$ satisfies the ISSC if the IFSs $\mathcal{I}_{C_{1}}$ and $\mathcal{I}_{C_{2}}$ satisfy the ISSC.
  2. (2)
    The IFS $\mathcal{P}_{C}$ satisfies the IOSC if both IFSs $\mathcal{I}_{C_{1}}$ and $\mathcal{I}_{C_{2}}$ satisfy the IOSC.
Proof.
  1. (1)
    Since $\mathcal{I}_{C_{1}}$ and $\mathcal{I}_{C_{2}}$ satisfy the ISSC, we have
    $f_{i}(A_{C_{1}})\cap f_{j}(A_{C_{1}})=\emptyset\quad\text{for all }i\neq j,\qquad f_{i}(A_{C_{1}})\cap C_{1}=\emptyset\quad\text{for all }1\leq i\leq N,$
    and
    $g_{k}(A_{C_{2}})\cap g_{l}(A_{C_{2}})=\emptyset\quad\text{for all }k\neq l,\qquad g_{k}(A_{C_{2}})\cap C_{2}=\emptyset\quad\text{for all }1\leq k\leq M.$
    This leads to
    $\Phi_{ij}(A_{C_{1}}\times A_{C_{2}})\cap\Phi_{kl}(A_{C_{1}}\times A_{C_{2}})=(f_{i}(A_{C_{1}})\times g_{j}(A_{C_{2}}))\cap(f_{k}(A_{C_{1}})\times g_{l}(A_{C_{2}}))=\emptyset,$
    for all $(ij)\neq(kl)$. Since $A_{\mathcal{P}_{C}}\subseteq A_{C_{1}}\times A_{C_{2}}$, it follows that
    $\Phi_{(ij)}(A_{\mathcal{P}_{C}})\cap\Phi_{(kl)}(A_{\mathcal{P}_{C}})=\emptyset\quad\text{for all }(ij)\neq(kl).$
    Furthermore,
    $\Phi_{ij}(A_{C_{1}}\times A_{C_{2}})\cap(C_{1}\times C_{2})=(f_{i}(A_{C_{1}})\times g_{j}(A_{C_{2}}))\cap(C_{1}\times C_{2})=\emptyset.$
    Therefore, the product IFS $\mathcal{P}_{C}$ satisfies the ISSC.
  2. (2)
    Since $\mathcal{I}_{C_{1}}$ and $\mathcal{I}_{C_{2}}$ satisfy the IOSC, there exist open sets $U\subset X$ and $V\subset Y$ such that
    $\bigcup_{i=1}^{N}f_{i}(U)\subset U,\quad f_{i}(U)\cap f_{j}(U)=\emptyset\quad\text{for all }i\neq j,\quad\text{and }U\subset\overline{C_{1}},$
    and
    $\bigcup_{k=1}^{M}g_{k}(V)\subset V,\quad g_{k}(V)\cap g_{l}(V)=\emptyset\quad\text{for all }k\neq l,\quad\text{and }V\subset\overline{C_{2}}.$
    This implies
    $\Phi_{(ij)}(U\times V)\cap\Phi_{(kl)}(U\times V)=(f_{i}(U)\times g_{j}(V))\cap(f_{k}(U)\times g_{l}(V))=\emptyset\quad\text{for all }(ij)\neq(kl).$
    Moreover,
    $\Phi_{(ij)}(U\times V)=f_{i}(U)\times g_{j}(V)\subset U\times V\quad\text{for all }(ij)\in\{1,\ldots,N\}\times\{1,\ldots,M\}.$
    Also, since $U\subset\overline{C_{1}}$ and $V\subset\overline{C_{2}}$, we have
    $U\times V\subset\overline{C_{1}}\times\overline{C_{2}}.$
    Define the open set $O:=U\times V\subset X\times Y$. Then
    $\bigcup_{(ij)}\Phi_{(ij)}(O)\subset O,\quad\Phi_{(ij)}(O)\cap\Phi_{(kl)}(O)=\emptyset\quad\text{for all }(ij)\neq(kl),\quad\text{and }O\subset\overline{C}.$
    Hence, the product IFS $\mathcal{P}_{C}$ satisfies the IOSC.
Let us see evidence of a connection between fractal measures of the IFSs $\mathcal{I}_{1}$, $\mathcal{I}_{2}$, and $\mathcal{P}_{C}.$
Theorem 7.3.
Let $\mathcal{I}_{1}=\{(X,d_{1});f_{1},f_{2},\ldots,f_{N},C_{1}\}$ and $\mathcal{I}_{2}=\{(Y,d_{2});g_{1},g_{2},\ldots,g_{M},C_{2}\}$ be two inhomogeneous iterated function systems (IFSs) with associated probability vectors $(p_{0},p_{1},\ldots,p_{N})$ and $(q_{0},q_{1},\ldots,q_{M})$, respectively. Define the inhomogeneous product IFS $\mathcal{P}_{C}:=\{X\times Y;\Phi_{(ij)},(C_{1},C_{2}),(ij)\in\{1,2,\ldots,N\}\times\{1,2,\ldots,M\}\}$ with $\Phi_{(ij)}(x,y)=(f_{i}(x),g_{j}(y))$ and probability vector
$\left\{r_{(ij)},\sum_{i=1}^{N}q_{0}p_{i},\sum_{j=1}^{M}p_{0}q_{j},p_{0}q_{0}~;~r_{(ij)}=p_{i}q_{j},1\leq i\leq N,1\leq j\leq M\right\}.$
Let $\nu_{1}$ and $\nu_{2}$ be Borel probability measures on $X$ and $Y$ with supports $C_{1}$ and $C_{2}$, respectively. Denote by $\mu_{\mathcal{I}_{1}}$ and $\mu_{\mathcal{I}_{2}}$ the inhomogeneous invariant measures associated with the IFSs $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$. Then the product measure $\mu_{\mathcal{I}_{1}}\times\mu_{\mathcal{I}_{2}}$ is an inhomogeneous-type measure, given by
$\mu_{\mathcal{I}_{1}}\times\mu_{\mathcal{I}_{2}}=\sum_{i=1,j=1}^{N,M}r_{ij}\mu_{\mathcal{P}_{C}}\circ\Phi_{ij}^{-1}+\left(\sum_{i=1}^{N}q_{0}p_{i}+\sum_{j=1}^{M}p_{0}q_{j}+p_{0}q_{0}\right)\nu,$
where $(\mu_{\mathcal{I}_{1}}\times\mu_{\mathcal{I}_{2}})(A\times B)=\mu_{\mathcal{I}_{1}}(A)\mu_{\mathcal{I}_{2}}(B)$ for Borel sets $A\subset X$, $B\subset Y$, and $\nu$ is a Borel probability measure on $X\times Y$.
Proof.
Since $r_{(ij)}=p_{i}q_{j}$, we have
$\mu_{\mathcal{I}_{1}}=\sum_{i=1}^{N}p_{i}\mu_{\mathcal{I}_{1}}\circ f_{i}^{-1}+p_{0}\nu_{1},\quad\mu_{\mathcal{I}_{2}}=\sum_{j=1}^{M}q_{j}\mu_{\mathcal{I}_{2}}\circ g_{j}^{-1}+q_{0}\nu_{2}.$
Then,
$\displaystyle\mu_{\mathcal{I}_{1}}\times\mu_{\mathcal{I}_{2}}$ $\displaystyle=\left(\sum_{i=1}^{N}p_{i}\mu_{\mathcal{I}_{1}}\circ f_{i}^{-1}+p_{0}\nu_{1}\right)\times\left(\sum_{j=1}^{M}q_{j}\mu_{\mathcal{I}_{2}}\circ g_{j}^{-1}+q_{0}\nu_{2}\right)$
$\displaystyle=\sum_{i=1,j=1}^{N,M}r_{(ij)}\mu_{\mathcal{I}_{1}}\circ f_{i}^{-1}\times\mu_{\mathcal{I}_{2}}\circ g_{j}^{-1}$
$\displaystyle\quad+\sum_{i=1}^{N}p_{i}q_{0}\mu_{\mathcal{I}_{1}}\circ f_{i}^{-1}\times\nu_{2}+\sum_{j=1}^{M}p_{0}q_{j}\nu_{1}\times\mu_{\mathcal{I}_{2}}\circ g_{j}^{-1}+p_{0}q_{0}\nu_{1}\times\nu_{2}$
$\displaystyle=\sum_{i=1,j=1}^{N,M}r_{(ij)}\left(\mu_{\mathcal{I}_{1}}\times\mu_{\mathcal{I}_{2}}\right)\circ\Phi_{(ij)}^{-1}$
$\displaystyle\quad+\left(\sum_{i=1}^{N}p_{i}q_{0}+\sum_{j=1}^{M}p_{0}q_{j}+p_{0}q_{0}\right)\nu,$
where
$\nu:=\frac{\sum_{i=1}^{N}p_{i}q_{0}\mu_{\mathcal{I}_{1}}\circ f_{i}^{-1}\times\nu_{2}+\sum_{j=1}^{M}p_{0}q_{j}\nu_{1}\times\mu_{\mathcal{I}_{2}}\circ g_{j}^{-1}+p_{0}q_{0}\nu_{1}\times\nu_{2}}{\sum_{i=1}^{N}p_{i}q_{0}+\sum_{j=1}^{M}p_{0}q_{j}+p_{0}q_{0}}$
is a Borel probability measure on $X\times Y$. ∎
8. Conclusion and future remarks
We have established a set of significant results on dimension theory and constructed new fractal sets using Falconer’s sub-self-similar concept. This work generalizes several existing results in the literature and, in the future, will be extended to more general fractal sets or attractors of dynamical systems. The fractal sets we have constructed have potential applications in the animation industry, computer-aided geometric design (CAGD), and graphic design. In future work, we plan to estimate the Assouad dimension of these ISSS sets and investigate the dimension spectrum.
Statements and Declarations
Data availability: Data sharing is not applicable to this article as no data sets were generated or analyzed during the current study.
Funding: The first author thanks IIIT Allahabad (Ministry of Education, India) for financial support through a Senior Research Fellowship. The second author acknowledges financial support from the SEED grant project of IIIT Allahabad.
Conflict of interest: We declare that we do not have any conflict of interest.
Author Contributions: All authors contributed equally to this manuscript.

### v2

@@
 1817, 24).
-We explore a range of examples of ISSS sets and elucidate a method to construct ISSS sets.
+We explore a range of ISSS sets and present a method for constructing them.
 We also investigate the upper and lower box dimensions of ISSS sets and discuss the continuity of the Hausdorff dimension.
-Key words and phrases:
-Self-similar sets, Inhomogeneous IFS, Hausdorff dimension, Box dimension, Open-set condition
 Key words and phrases:
@@
 Priyadarshi [34] estimated the lower bound on the Hausdorff dimension of sets of complex continued fractions using operator-theoretic techniques.
-Chandra and Abbas [8] studied the fractal dimension of fractal functions using function space techniques.
+Chandra and Abbas [8] studied the fractal dimension of fractal functions using function space technique.
 Falconer [12] also explored fractal dimensions using potential-theoretic techniques and the mass distribution principle.
-Hochman [19] investigated the Hausdorff dimension of self-similar sets and measures on the real line using entropy techniques and provided an inverse theorem for the entropy of convolution.
+Hochman [19] investigated the Hausdorff dimension of self-similar sets and measures on the real line using entropy technique and provided an inverse theorem for the entropy of convolution.
 Shmerkin [40] extended this work by studying the $L^{q}$-dimension of dynamically driven self-similar measures.
@@
 In Section 3, we present examples of ISSS sets.
-Section 4 discusses one of the methods for constructing ISSS sets, while Sections 5 and 6 are devoted to estimating the fractal dimension of these sets and discussing the continuity of the Hausdorff dimension for ISSS sets.
+Section 4 discusses a method for constructing ISSS sets, while Sections 5 and 6 are devoted to estimating the fractal dimension of these sets and examining the continuity of the Hausdorff dimension for ISSS sets, respectively.
 In Section 7, we study the self-similar structure on the product of inhomogeneous IFS.
-Finally, Section 8 concludes the paper with remarks and outlines potential future directions for research.
+Finally, Section 8 concludes the paper with remarks and outlines for potential future research directions.
 2. Preliminaries
@@
 2.1. Code Space
-Let $I^{\infty}$ denote the set of all infinite sequences in $\{1,\ldots,N\}$, that is,
+Let $I^{\infty}$ denotes the set of all infinite sequences in $\{1,\ldots,N\}$, that is,
 $I^{\infty}:=\left\{\{\omega_{k}\}_{k=1}^{\infty}\,\middle|\,\omega_{k}\in\{1,\ldots,N\}\right\}.$
-An element of $I^{\infty}$ is written as $\omega:=\omega_{1}\omega_{2}\dots\omega_{k}\omega_{k+1}\dots$, and the $k$-th term of the sequence is denoted by $\omega_{k}$.
+We write an element of $I^{\infty}$ as $\omega:=\omega_{1}\omega_{2}\dots\omega_{k}\omega_{k+1}\dots$, and the $k$-th term of the sequence is denoted by $\omega_{k}$.
 For any subset $S\subseteq I^{\infty}$ and a positive integer $n$, define
-$S^{n}:=\left\{\omega_{|_{n}}=\omega_{1}\dots\omega_{n}\,\middle|\,\omega=\omega_{1}\omega_{2}\dots\omega_{k}\dots\in S\right\},\quad\text{and}\quad S^{*}:=\bigcup_{n}S^{n}.$
+$S^{n}:=\left\{\omega_{|_{n}}=\omega_{1}\dots\omega_{n}\,\middle|\,\exists~\omega=\omega_{1}\omega_{2}\dots\omega_{k}\dots\in S\right\},\quad\text{and}\quad S^{*}:=\bigcup_{n}S^{n}.$
 Suppose $(s_{1},s_{2},\ldots,s_{N})$ is a collection of real numbers such that $0<s_{i}<1$ for all $1\leq i\leq N$. A metric $d_{\infty}$ on $I^{\infty}$ is defined by
@@
 which associates each infinite sequence $\omega$ with a unique point in the attractor $A$.
-A compact subset $K\subseteq I^{\infty}$ is said to be closed under left shift if, for every $\omega=\omega_{1}\omega_{2}\dots\in K$, the sequence $\omega^{\prime}:=\omega_{2}\omega_{3}\dots$ also belongs to $K$.
+A compact subset $K\subseteq I^{\infty}$ is said to be closed under left shift if for every $\omega=\omega_{1}\omega_{2}\dots\in K$, the sequence $\omega^{\prime}:=\omega_{2}\omega_{3}\dots$ also belongs to $K$.
 Falconer characterizes sub-self-similar sets and establishes a natural correspondence between such sets and subsets of the code space.
@@
 Unlike inhomogeneous self-similar sets, ISSS sets are not unique.
-For example, consider the IFS $\mathcal{I}=\{f_{1}(x)=\frac{x}{2},f_{2}(x)=\frac{x}{2}+\frac{1}{2}\}$ with condensation set $C$.
+For example, consider the IFS $\mathcal{I}=\{f_{1}(x)=\frac{x}{2},f_{2}(x)=\frac{x}{2}+\frac{1}{2}\}$ with a condensation set $C$.
 Both the set $\{0,1\}$ and the set of all dyadic numbers in $[0,1]$ are ISSS sets corresponding to $\mathcal{I}$ and $C$.
+In particular, if we consider $C=\{\frac{1}{2}\},$ then set $\{\frac{1}{2^{n}}:n\geq 1\}\cup\{1+\frac{1}{2^{n}}:n\geq 1\}$ is an ISSS with condensation set $C$ set but not an SSS set.
 Remark 3.2.
@@
 In [31], McClure and Vallin provide examples of non-sub-self-similar sets, such as $E=\{0,1,\frac{1}{2},\frac{1}{3},\ldots\}$, which is not a sub-self-similar set.
-According to the previous remark, there exists a condensation set $C$ such that $E$ becomes an ISSS set.
+As noted previously, there exists a condensation set $C$ such that $E$ becomes an ISSS set.
 Furthermore, for any compact set $E$, the minimal condensation set $C=\overline{E\setminus\bigcup_{i=1}^{N}f_{i}(E)}$ ensures that $E$ is an ISSS set.
@@
 Let $E$ be a sub-self-similar set associated with the IFS $\mathcal{I}$ such that $\Theta(S)=E$, and let $C\subset X$ be a non-empty compact set. Define
-$O_{S}=\bigcup_{{\omega}\in S^{*}}f_{{\omega}}(C).$
+$O_{S}:=\bigcup_{{\omega}\in S^{*}}f_{{\omega}}(C).$
 Then $E\cup O_{S}$ is an ISSS set corresponding to the IFS $\mathcal{I}$ with condensation set $C$.
@@
 Corollary 5.2.
-Let $E\cup O_{S}$ be an ISSS set with SSS set $E$ and condensation set $C.$ Then $E\cup O_{S}=\max\{s,\dim_{H}C\}$.
+Let $E\cup O_{S}$ be an ISSS set with SSS set $E$ and condensation set $C.$ Then $\dim_{H}(E\cup O_{S})=\max\{s,\dim_{H}C\}$.
 Proof.
@@
 Proof.
-The function $\tau(s)=\lim_{k\to\infty}\bigg{(}\sum_{\omega\in S^{k}}\rho_{\omega}^{s}\bigg{)}^{1/k}$ defined in Theorem (5.1) is strictly decreasing continuous function with $0\leq\tau(s)<\infty$ and $\sum_{\omega\in S^{k}}\rho_{\omega}^{s}\geq\tau(s)^{k}$ for all $k=1,2,\ldots$
+The function $\tau(s)=\lim_{k\to\infty}\bigg(\sum_{\omega\in S^{k}}\rho_{\omega}^{s}\bigg)^{1/k}$ defined in Theorem 5.1 is strictly decreasing continuous function with $0\leq\tau(s)<\infty$ and $\sum_{\omega\in S^{k}}\rho_{\omega}^{s}\geq\tau(s)^{k}$ for all $k=1,2,\ldots$
 Also observe that $\sum_{\omega\in S^{*}}\rho_{\omega}^{t}=\sum_{k=1}^{\infty}\sum_{\omega\in S^{k}}\rho_{\omega}^{t}$ converges if $\tau(s)<1$ and diverges if $\tau(s)>1.$ Hence for $t>s$$\sum_{\omega\in S^{*}}\rho_{\omega}^{t}=\sum_{k=1}^{\infty}\sum_{\omega\in S^{k}}\rho_{\omega}^{t}=m_{t}<\infty,$ completing the assertion.
@@
 $\max\{\overline{\dim}_{B}E,\overline{\dim}_{B}C\}\leq\overline{\dim}_{B}E_{C}\leq\max\{\overline{\dim}_{B}E,\overline{\dim}_{B}O_{S}\}.$
-By Theorem (5.1) $\overline{\dim}_{B}E\leq s,$ it is sufficient to show that $\overline{\dim}_{B}O_{S}\leq\max\{s,\overline{\dim}_{B}C\}.$ Let $\max\{s,\overline{\dim}_{B}C\}<t,$ then by the definition of the box-dimension, there exists some constant $c_{t}$ such that
+By Theorem 5.1 $\overline{\dim}_{B}E\leq s,$ it is sufficient to show that $\overline{\dim}_{B}O_{S}\leq\max\{s,\overline{\dim}_{B}C\}.$ Let $\max\{s,\overline{\dim}_{B}C\}<t,$ then by the definition of the box-dimension, there exists some constant $c_{t}$ such that
 $N_{\delta}(C)\leq c_{t}\delta^{-t},$
@@
 $\leq(c_{t}+c_{t}m_{t}+m_{t}{\rho}^{-t}_{\min}N_{1}(X))\delta^{-t}.$
-The third line comes from Lemma (5.6). The sixth line comes as $X$ is a compact space and $\delta/\rho_{\omega}>1,$ the last line arrived by Lemma (5.4). This completes the proof. ∎
+The third line comes from Lemma 5.6. The sixth line follows because $X$ is a compact space and $\delta/\rho_{\omega}>1$, as established by Lemma 5.4. This completes the proof. ∎
 5.3. Lower box dimension
@@
 $N_{\delta_{0}}(C)\leq\delta_{0}^{-t}.$
-By Lemma (5.4), for $t>s$, the cardinality of the $\delta$-stopping set $S$ is given by:
+By Lemma 5.4, for $t>s$, the cardinality of the $\delta$-stopping set $S$ is given by:
 $|S(\delta)|\leq m_{t}\rho_{\min}^{-t}\delta^{-t}.$
@@
 Theorem 5.10.
-For all $t>\max\{s,\underline{\dim}_{B}C\}$, we have:
+We have:
 $\underline{\dim}_{B}(E\cup O_{S})\leq\max\{s,\underline{\dim}_{B}C\}.$
@@
 Let $t>\max\{s,\underline{\dim}_{B}C\}$.
-Then, by Lemma (5.8), we have $P_{t}(C)\leq\underline{\dim}_{B}C<t$.
-Thus, by Lemma (5.9), for all $\epsilon\in(0,1-P_{t}(C))$, there exists $\delta\in(0,\epsilon)$ such that:
+Then, by Lemma 5.8, we have $P_{t}(C)\leq\underline{\dim}_{B}C<t$.
+Thus, by Lemma 5.9, for all $\epsilon\in(0,1-P_{t}(C))$, there exists $\delta\in(0,\epsilon)$ such that:
 $P_{t}(C)-\epsilon<P_{t,\delta}(C)<P_{t}(C)+\epsilon.$
@@
 Proof.
-Let $I_{k}=\{f_{\omega}:\omega\in S^{k}\}$. For any $x\in F=E\cup O_{S}$, we consider two cases:
+Let $S^{k}:=\{\omega=\omega_{1}\omega_{2}\ldots\omega_{k}:\exists\omega_{1}\ldots\omega_{n}\ldots\in S\}.$ Define IFSs $I_{k}:=\{f_{\omega}:\omega\in S^{k}\}$ with corresponding ISSS set $F_{k}$ for each $k\in\mathbb{N}$.
+Then, For any $x\in F=E\cup O_{S}$, we consider two cases:
 Case 1: If $x\in E$, then there exists a string $\omega=\omega_{1}\omega_{2}\ldots\in S$ such that
 $x=\lim_{n\to\infty}f_{\omega_{1}}\circ f_{\omega_{2}}\circ\ldots\circ f_{\omega_{n}}(z).$
-For each $k\in\mathbb{N}$, define $\omega_{j}|_{k}^{k}:=\omega_{jk+1}\omega_{jk+2}\ldots\omega_{jk+k}$ for $j=0,1,2,\ldots$. Then $\omega_{j}|_{k}^{k}\in S^{k}$, and we have
-$x=\lim_{j\to\infty}f_{\omega_{0}|_{k}^{k}}\circ f_{\omega_{1}|_{k}^{k}}\circ\ldots\circ f_{\omega_{j}|_{k}^{k}}(z),$
+For each $k\in\mathbb{N}$, we rewrite $\omega$ as $\omega:=(\omega_{1}\omega_{2}\ldots\omega_{k})|_{0}(\omega_{k+1}\omega_{k+2}\ldots\omega_{2k})|_{1}\ldots(\omega_{jk+1}\omega_{jk+2}\ldots\\ \omega_{(j+1)k})|_{j}\ldots\in S.$ Since, $S$ is closed under left shift, it implies that each block $\omega_{j}:=\omega_{jk+1}\ldots\omega_{(j+1)k}\in S^{k}$ for each $j=0,1,\ldots$, and we have
+$x=\lim_{j\to\infty}f_{\omega_{0}}\circ f_{\omega_{1}}\circ\ldots\circ f_{\omega_{j}}(z),$
 which implies that $x\in\bigcup_{k=1}^{\infty}F_{k}$.
@@
 Now suppose $x\in F_{k}$ for some $k$.
-Then there exists a sequence $\omega_{0}|_{k}^{k},\omega_{1}|_{k}^{k},\ldots,\omega_{j}|_{k}^{k}\in S^{k}$ such that
-$x=\lim_{j\to\infty}f_{\omega_{0}|_{k}^{k}}\circ f_{\omega_{1}|_{k}^{k}}\circ\ldots\circ f_{\omega_{j}|_{k}^{k}}(z).$
-Using the fact that $\Theta(S)=E$ and that the concatenation $\omega=\omega_{0}|_{k}^{k}\omega_{1}|_{k}^{k}\ldots\omega_{j}|_{k}^{k}\in S$, we conclude that $x\in F$.
+Then there exists a sequence $\{\omega_{j}\}_{j=1}^{\infty}$ in $S^{k}$ such that
+$x=\lim_{j\to\infty}f_{\omega_{1}}\circ f_{\omega_{2}}\circ\ldots\circ f_{\omega_{j}}(z).$
+Using the fact that $\Theta(S)=E$ and that the concatenation $\omega=\omega_{1}\omega_{2}\ldots\omega_{j}\ldots\in S$, we conclude that $x\in F$.
 Thus, we obtain $F=\bigcup_{k=1}^{\infty}F_{k}$. ∎
+Note 6.2.
+It is worth noting that, if the underlying IFS satisfies the OSC, then the IFS $I_{k}$ also satisfy the OSC for each $k\in\mathbb{N},$ and $\{E_{k}\}_{{k=1}^{\infty}}$ is a monotonic increasing sequence.
+If $x\in E_{k},$ then there exists a sequence $\{\omega_{j}\}_{j=1}^{\infty}$ in $S^{k}$ such that $\lim_{j\to\infty}f_{\omega_{1}}\circ f_{\omega_{2}}\circ\cdots\circ f_{\omega_{j}}(z)=x.$ By the fact that $x\in F$ and $\Theta(S)=E,$ the concatenation $\Omega:=\omega_{1}\ldots\omega_{j}\ldots\in S.$ Since, $\omega_{j}\in S^{k},$ by our construction there exist $\omega^{j}\in S$ such that $\omega^{j}|_{k}=\omega_{j}:=\omega^{j}_{1}\ldots\omega^{j}_{k}$ for each $j\in\mathbb{N}.$ Now, we rewrite $\Omega=(\omega^{1}_{1}\ldots\omega^{1}_{k}\omega^{2}_{1})(\omega^{2}_{2}\ldots\omega^{2}_{k}\omega^{3}_{1}\omega^{3}_{2})\ldots$, then clearly each block of length $k+1$ and $S^{k+1}.$ This implies that $x\in E_{k+1}.$
 Lemma 6.3.
@@
 Corollary 6.4.
-Let $F=E\cup O_{S}$ be the ISSS set corresponding to the IFS $\{f_{i}\}{i=1}^{N}.$ Then there exists a sequence of IFSs $\{I_{k}\}$ with ISSS sets $\{F_{k}\}$ and under inhomogeneous open set condition (IOSC) $\dim_{H}E=\lim_{k\to\infty}\dim_{H}E_{k}=\lim_{k\to\infty}=\lim_{k\to\infty}s_{k},$ where $s_{k}$ is the Hausdorff dimension of $F_{k}$ defined in Lemma 6.2.
+Let $F=E\cup O_{S}$ be the ISSS set corresponding to the IFS $\{f_{i}\}_{i=1}^{N}.$ Then, there exists a sequence of IFSs $\{I_{k}\}_{k=1}^{\infty}$ with ISSS sets $\{F_{k}\}$ and under inhomogeneous open set condition (IOSC) $\dim_{H}E=\lim_{k\to\infty}\dim_{H}E_{k}=\lim_{k\to\infty}s_{k},$ where $s_{k}$ is the Hausdorff dimension of $F_{k}$ defined in Lemma 6.3.
 Proof.
-By Theorem 6.1, there exists a sequence of IFSs $\{I_{k}\}$ with ISSS sets $F_{k}$ such that $F=\bigcup_{k=1}^{\infty}F_{k}.$ It is easy to observe that if $\{f_{i}\}_{i=1}^{N}$ satisfies the IOSC, then each $I_{k}$ also satisfies the IOSC.
+By Theorem 6.1, there exists a sequence of IFSs $\{I_{k}\}_{k=1}^{\infty}$ with ISSS sets $F_{k}$ such that $F=\bigcup_{k=1}^{\infty}F_{k}.$ It is easy to observe that if $\{f_{i}\}_{i=1}^{N}$ satisfies the IOSC, then each $I_{k}$ also satisfies the IOSC.
 If $s_{k}$ represents the Hausdorff dimension of $E_{k}$, which satisfies
@@
 then the Hausdorff dimension of $F_{k}$ is $\max\{s_{k},\dim_{H}(C)\}$.
-By Lemma (6.2), we have
+By Lemma 6.3, we have
 $\dim_{H}E=\lim_{k\to\infty}\dim_{H}E_{k}=\lim_{k\to\infty}s_{k},$